Write the isotope needed, by identifying A, Z and X to balance the following nuclear fission reaction:

235 1 143 A 192 U+ 0n→ 55 Cs+ Z X +3 0 n

Express your answer as an isotope using prescripts.



1
Expert's answer
2021-02-26T06:01:47-0500

"\frac{235}{92}U + \frac{1}{0}n = \frac{137}{55}Cs + \frac{96}{37}Rb + 3 \frac{1}{0}n"